OpenStudy (anonymous):

Choose the correct description of the graph of the inequality x - 4 > 5x - Open circle on -1, shading to the left. Closed circle on -1, shading to the left. Open circle on -1, shading to the right. Closed circle on -1, shading to the right.

OpenStudy (anonymous):

*asking for trouble* what do you think?

OpenStudy (anonymous):

3rd one?

OpenStudy (anonymous):

did u get \[x\leq -1\;?\]

OpenStudy (anonymous):

no im confused.

OpenStudy (anonymous):

okay wait...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the 2nd one ?

OpenStudy (goformit100):

2nd one is correct

OpenStudy (anonymous):

\[x-4\geq 5x\\-4\geq4x\\-1\geq x\\x\leq-1\]

OpenStudy (anonymous):

Thank yall :)